### v1

# Automatic Mapping of AutomationML Files to Ontologies for Graph Queries and Validation

Abstract
AutomationML has seen widespread adoption as an open data exchange format in the automation domain. It is an open and vendor neutral standard based on the extensible markup language XML. However, AutomationML extends XML with additional semantics, that limit the applicability of common XML-tools for applications like querying or data validation. This article provides practitioners with 1) an up-to-date ontology of the concepts in the AutomationML-standard, as well as 2) a declarative mapping to automatically transform any AutomationML model into RDF triples. Together, these artifacts allow practitioners an easy integration of AutomationML information into industrial knowledge graphs. A study on examples from the automation domain concludes that transforming AutomationML to OWL opens up new powerful ways for querying and validation that are impossible without transformation.
This research is also funded by the Federal Ministry for Economic Affairs and Climate Action (BMWK) on the basis of a decision by the German Bundestag.
Keywords: AutomationML, Ontologies, Graph Queries, Validation, Semantic Web, Industry 4.0
1 Introduction
AutomationML is an open and vendor neutral solution for data exchange in the domain of automation engineering. Since its first publication in 2008 [1], it has seen widespread adoption in both the scientific community as well as industry, where it is viewed as a central building block for digital twins [2].
AutomationML is designed to represent complex relationships within engineering projects. It is used to model and exchange data related to the design, implementation, and maintenance of automation systems. The AutomationML data model takes the form of a deeply interconnected graph. However, for easier exchange with other engineering tools, AutomationML uses XML as its underlying format for serialization, which follows a tree structure. In the serialization, references between elements of the XML-tree are realized through reference paths and Identifiers.
In practice, this representation mechanism offers multiple drawbacks. First, common tools to query and validate XML-Files (like XSD or xQuery) were designed to operate on the hierarchical structure of the data format. They therefore have very rudimentary support for operations involving the additional connections defined by AutomationML. Second, the data lacks a formal representation which would allow knowledge inference and further analytical applications[3]. Third, the XML-Schema of AutomationML does not allow for native alignment to other data sources, i.e. it is limited to the vocabulary that AutomationML has to offer. However, this alignment is a pillar of the digital twin[2], making it especially necessary in the engineering domain [4].
In terms of formally representing engineering knowledge, technologies like the Resource Description Framework (RDF) and the Web Ontology Language (OWL), are widely used, as they provide a standardized framework for expressing relationships and semantics [5]. These technologies have been applied to create meta models of cyber-physical systems [6][7], which allows for the integration of heterogeneous data and resolving semantic heterogeneity [4]. RDF and OWL further allow the use of technologies from the Semantic Web Stack like SPARQL and SHACL, which further enable efficient querying and validation of the created knowledge graphs.
The transformation of AutomationML into an ontology-compliant knowledge graph has been in the focus of the research community for a while[8, 9, 10, 3, 11], usually with a focus on data integration.
However, each of these proposed AutomationML ontologies exhibits significant shortcomings in their scope, design methodology, or compliance with current AutomationML specifications. Additionally, there is no freely available tool or mapping definition to automatically transform an AutomationML file into an RDF graph, further limiting widespread adoption.
Thus, this article aims to close these gaps by providing practitioners with two artifacts to be reused in future works:
1.
An up to date AutomationML Ontology, with full coverage of the standards vocabulary.
2.
A declarative mapping that offers practitioners a straightforward transformation of AutomationML files to RDF triples.
Using these artifacts, it is possible to automatically translate an AutomationML file into an ontology. This translation process enables new use cases for querying and validating AutomationML files, which are demonstrated on an example system.
The remainder of the article is structured as follows: Section 2 gives more information on the AutomationML format, how it relates to graph theory and on the tools from the Semantic Web Stack. Section 3 describes previous attempts at an AutomationML ontology, while Section 4 discusses the necessary changes and additions to the AutomationML ontology. Section 5 describes the design decisions of mapping AutomationML files to RDF. The mapping’s correctness and flexibility are then validated on multiple AutomationML files in Section 6. Here, the AutomationML files are used for various query and data validation scenarios. The use cases show that transforming AutomationML to RDF allows for more complex graph queries and validation rules in comparison to typical XML-based approaches.
2 Background
2.1 AutomationML
AutomationML (Automation Markup Language) is an open, XML-based standard for the exchange of engineering data. It builds on established and standardized data exchange formats like Computer Aided Engineering Exchange (CAEX) for plant topology, COLLADA for geometry and kinematics, and PLCopenXML for control logic. It is commonly used in both the scientific community and industry[1, 12] and has subsequently been standardized in IEC 62714[13].
Recently, a new normative document - the IEC 62714-2[14] - was released, which expands the scope of the AutomationML standard by introducing semantics libraries. Including the newly added AttributeTypeLibrary, the CAEX submodel of AutomationML now consists of five main pillars (see Fig. 1)[15]:
The InstanceHierarchy is a hierarchy of InternalElements and contains the data of a specific project.
The SystemUnitClassLibrary provides SystemUnitClasses, which describe types of physical plant objects, e.g. a specific type of robot or valve.
The RoleClassLibrary provides RoleClasses, which define abstract roles that a physical or logical plant objects may fill. While SystemUnitClasses usually describe some manufacturer specific systems, RoleClasses model the abstract functions of components.
The InterfaceClassLibrary provides InterfaceClasses which define a type of interface (e.g. a flange or a signal interface). These interfaces are used to describe relationships between InternalElements (IE) and SystemUnitClasses (SUC) through InternalLinks.
The AttributeTypeLibrary defines a predefined data structure of reusable AttributeTypes and can be used to model attributes with values, units and descriptions. It allows for the definition of proprietary or standardized vocabularies of attributes, and creates a basis for the semantic interpretability of attributes.
The content of an AutomationML file is structured in the shape of an XML-hierarchy, in which each element has a specific path. Additionally, AutomationML makes intensive use of references and links between different elements on different branches on this tree. These references between elements of the XML-tree are realized through reference paths. Using this mechanism, classes and roles can be linked to instances of this class. It further allows for the definition of interfaces, which function as connection points between different AutomationML elements.
2.2 AutomationML and Graph Theory
In mathematical terms, a graph is defined as an ordered pair:
where:
is a set of vertices (nodes), and
is a set of edges, where each edge is a pair of distinct vertices:
Graphs can take on various shapes and possess different properties, depending on how their vertices and edges are arranged.
XML-based files (like AutomationML) can be considered as a specific subtype of graphs. Here, the XML-Elements are considered as the nodes, whereas the hierarchical organization of elements forms the edges. If viewed in this way, XML files naturally take the shape of a Tree (s. left side of Figure 2). Trees are a specific type of acyclic, connected graph:
where:
Each node has a unique path from the root, ensuring a clear hierarchical organization of data.
Every edge is directed, so for each edge , is considered the parent and the child.
XML’s hierarchical nature and the unique path from the root to each element ensure that the data can be traversed, queried, and manipulated as a tree, making it an ideal format for structured data representation. XML-based tools for querying and validation heavily rely on this hierarchical parent / child structure.
However, as stated in Section 2.1, AutomationML also allows additional references between nodes besides their hierarchical position in the XML file. Once the ReferencePaths and InternalLinks are considered, the AutomationML model no longer satisfies the property of having a unique path between every pair of nodes. An example of such a cyclical AutomationML graph can be seen on the right side of Figure 2. It is therefore no longer a Tree as defined by Definition 2
The data instead forms a labeled directed multigraph. It is defined as a triple
where:
is a set of vertices (or nodes),
is a multiset of ordered pairs of directed edges. Each edge is represented as , where , with as the source and as the target, and
is a function that assigns a label to each edge in , mapping , where is a set of labels.
In this type of graph, multiple edges can exist between the same nodes. Edges are described further by their labels, that define which type of edge (e.g. RefBaseClass, InternalLink etc.) exists between the nodes.
For the use of XML tools on AutomationML files, this mismatch between XML’s tree structure and AutomationML’s multigraph structure results in some restrictions.
XML tools for queries (like xPath and xQuery) as well as for validation (like XSD and Schematron) were designed to handle hierarchical data in the shape of Trees (s. Equation 2). They therefore only offer limited support for connections that are not part of the XML tree structure. For AutomationML, this includes any connection that involves referencing another element via either reference path or ID, i.e. any class assignment or InternalLink between elements.
For data in the shape of labeled directed multigraphs, other data formats and technologies exist, that are able to natively deal with the data’s cyclical and interconnected nature. Semantic Web Technologies are commonly used for this task and will be explained in more detail in the following Section.
2.3 Semantic Web Technologies
Semantic Web Technologies are a framework of technologies designed to enable machines to meaningfully interpret and process data. The Resource Description Framework (RDF), and its Schema Extension (RDFS) and the Web Ontology Language (OWL) are related technologies used in the context of the Semantic Web Technologies. They are often used together to represent and reason about knowledge in a machine-readable format.
RDF is a framework for representing information about resources on the web. It provides a standardized means to describe resources using subject-predicate-object triples. Each triple represents a statement or a fact about a resource. Mathematically speaking, an RDF-Graph takes the shape of a labeled and directed multigraph, as it allows different types of directed edges.
The Web Ontology Language OWL, on the other hand, is a language for defining ontologies. An ontology is a formal description of concepts, relationships, and properties within a certain domain[16]. OWL allows for more expressive and complex modeling of relationships between entities than RDF: While RDF provides a foundation for expressing relationships, OWL enhances the expressiveness and formalism of these relationships, enabling the creation of ontologies for more advanced knowledge representations. In combination with a reasoner, OWL enables the inference of additional knowledge.
The RDF Mapping Language (RML) is a mapping language defined to express customized mapping rules from heterogeneous data structures and serializations to the RDF data model[17]. It has seen widespread use since its inception and can be applied to structured data sources like CSV, JSON or XML-based formats like AutomationML. Numerous platform independent and open source tools are available to automatically convert data sources to RDF-Triples, given a suitable mapping.
SPARQL is a query language and protocol designed for querying and manipulating data stored in RDF format. It can be used to express patterns and conditions in order to retrieve, construct, analyze, or modify information stored in RDF graphs[18].
The Shapes Constraint Language (SHACL) is a language used for describing and validating the structure and content of data in RDF graphs. SHACL provides a powerful mechanism for defining and enforcing data validation rules in RDF graphs, helping ensure data quality and consistency in semantic web applications[19].
3 Requirements and State of the Art
In the past, numerous attempts have been made to create a formalized AutomationML ontology. The following Section first outlines the requirements that a usable AutomationML ontology needs to fulfill. Afterwards it gives an overview of related works and previous AutomationML ontologies, before comparing previous ontologies to the requirements.
3.1 Requirements
The resulting knowledge graph should be both highly expressive as well as easy to use for people that are already familiar with AutomationML’s terminology. This results in a number of requirements, that influence later design decisions. For instance, in order to make the ontology easy to understand and use for AutomationML-practitioners, new terminology should be avoided, resulting in Requirement R1. Since queries as well as validation use cases make use of the entire width of information that is stored in an AutomationML file, the ontology should likewise cover the entire vocabulary that CAEX has to offer (R2). However, the CAEX vocabulary within AutomationML is subject to contextual interpretation based on hierarchical positioning and inter-element relationships. For example the RefBaseSystemUnitClass-Property can be used to define the SystemUnitClass of an InternalElement. Used in a different context however, it can also describe a Mirror/Master relationship between two InternalElements (s. Sec. 5.7). Recognizing these contextual semantics and making them explicitly available are important for both accurate modeling as well as usability of the resulting knowledge graph (R3).
AutomationML is an actively maintained and developed data exchange format. The ontology should therefore consider the most recent version of the AutomationML specification (R4).
Most applications of Semantic Web Technologies make heavy use of OWL terminology concepts like class/property hierarchies and property characteristics. The ontology should therefore enrich and align AutomationML concepts with the terminology used in OWL (R5).
Manually translating an AutomationML file to OWL involves extensive effort and can introduce ambiguities in the mapping process. In order to allow practitioners to actually use the ontology beyond the scope of prototypical implementations, a reusable method to automatically translate AutomationML files to ontologies is needed (R6).
- R1.
Class/Property names should follow the AutomationML terminology.
- R2.
The ontology should cover the entire CAEX vocabulary.
- R3.
The ontology should allow explicit modeling of context-dependent AutomationML concepts.
- R4.
The ontology should follow the current AutomationML specification.
- R5.
Enrich and align concepts from AutomationML with OWL concepts.
- R6.
A mapping/translation approach for the automatic conversion from AutomationML files to the ontology must be available.
3.2 Related Works: Mapping and Applications
With regard to mapping AutomationML files to OWL, Hua et al. [20] introduce AutomationML concept models for representing OWL complex classes in AutomationML. Additionally, the authors present algorithms for the bidirectional translation between OWL complex classes and their corresponding AutomationML concepts. Their focus, however, is on leveraging the semantics of the AutomationML concept models, not on providing practitioners with a straightforward method of translation between the two representations.
Sabou et al. [21] present the AutomationML-Analyzer, which is a tool that showcases the functionalities made possible by the combination of AutomationML files with semantic web technologies. The authors use the aforementioned AutomationML Ontology presented in [22]. The mapping is performed using a two-step process, in which the AutomationML-files are first converted to Ecore, a domain model of the Eclipse Modeling Framework, and then converted from Ecore to OWL. This conversion procedure is identified as less than ideal by the authors themselves, as they propose a direct mapping as a potential future work. With regards to querying AutomationML-files, Wimmer et al. [23] define an AutomationML-specific query language. While it expands upon the capabilities of other XML-based Query languages, it similarly uses the Ecore-Mapping.
The formal representation of the knowledge in an AutomationML file has been used as the basis for numerous applications over the past years. For example, Abele et al. [8] transformed CAEX-models of production plants to OWL and subsequently used the resulting formal representation to perform a design validation using semantic web reasoning.
Grangel-Gonzales et al. [10] used knowledge from AutomationML files to semantically describe cyber-physical systems from different design perspectives. Gonçalves et al. [11] use the engineering knowledge contained in an AutomationML file with an ontology for sensor fusion to deal with the occurrence of data divergence.
3.3 Prior AutomationML-Ontologies
In the past, multiple attempts at an AutomationML ontology have been published. Abele et al.[8] cover the most important aspects of the CAEX vocabulary, and even some basic context-dependent concepts like MirrorObjects. However, neither of those are complete as lots of additional information about the entities is not covered. An automated mapping approach using Apache Jena is mentioned, but unavailable.
Grangel-Gonzales et al. [9] focus on the Attribute structure of AutomationML and are therefore lacking coverage of the remaining CAEX vocabulary. AutomationML files are mapped automatically using a XSLT-based tool. Due to the limited coverage of the model, the mapping method is not viable for large scale projects.
Kovalenko et al. [22] focus on the description of the CAEX vocabulary, but without covering the Attributes. An automated mapping approach using Apache Jena is mentioned, but unavailable.
Kovalenko et al. later present an Ontology of the CAEX-submodel of the AutomationML-standard [3]. This version is essentially a combination of the previous works from [9] and [22]. The ontology provides fairly good coverage of the basic CAEX concepts, with some notable exceptions like constraints. The context dependent concepts (like e.g. Mirror- or MappingObjects) remain missing.
Since the release of all considered ontologies, the AutomationML standard underwent continuous development and has since been released as Version 3.0 in 2022[14]. All ontologies were released before this date and can therefore not fulfill Requirement R4. Because of this, they lack concepts for the semantic AttributeLibraries introduced in [14].
Previous Ontology | [8] | [22] | [9] | [3] |
R1: Terminology | ||||
R2: Complete Coverage | ||||
R3: Context-dependent Entities | ||||
R4: Current AutomationML Version | ||||
R5: Enrich Model with OWL concepts | ||||
R6: Automated Mapping Approach |
As Table 2 shows, none of the considered Ontologies fulfill all requirements. However, the ontology presented by Kovalenko et al. [3] achieves good coverage of the basic CAEX concepts. Since the new AutomationML version[14] introduces significant changes to the AutomationML vocabulary as well as the XML serialization, the decision was made to develop a new ontology. Since many the basic CAEX concepts did not change, and [3] achieved good coverage in that area, their ontology was used as a starting point.
4 AutomationML Ontology
As described in Section 3.2, a new AutomationML ontology was designed that aims to to fulfill the Requirements outlined in Section 3.1. The section first gives an overview of the CAEX vocabulary that could be reused (R2), before describing the concepts that had to be added to fulfill Requirements R3 and R4.
4.1 Reused entities from previous AutomationML Ontology
The AutomationML Ontology from Kovalenko et al. [3] already offered strong coverage of the basic CAEX concepts. Because of this, most of the common concepts could be reused in the new ontology (s. Table 3). Many additional concepts are present in the ontology, but needlessly differ from the AutomationML terminology. These have been renamed to align them with Requirement R1.
AutomationML Concept | Ontology Class |
CAEXFile | :CAEXFile |
InstanceHierarchy | :InstanceHierarchy |
SystemUnitClassLib | :SystemUnitClassLib |
RoleClassLib | :RoleClassLib |
InterfaceClassLib | :InterfaceClassLib |
InternalElement | :InternalElement |
SystemUnitClass | :SystemUnitClass |
RoleClass | :RoleClass |
InterfaceClass | :InterfaceClass |
ExternalInterface | :ExternalInterface |
Attribute | :Attribute |
ExternalReference | :ExternalReference |
InternalLink | :InternalLink |
AutomationML Concept | Object Property |
InternalStructure | :hasPart, :hasInterface, :hasAttribute |
InternalLink | :hasRefPartnerSideA, :hasRefPartnerSideB |
RoleRequirements | :hasRoleRequirement |
SupportedRoleClass | :hasSupportedRoleClass |
4.2 Additional CAEX Vocabulary
In part due to continued development of the AutomationML-Standard, some parts of the CAEX vocabulary had to be added to fulfill Requirement R2. This mostly refers to concepts needed to define the semantics of attributes. By combining these new concepts with the reused concepts from Table 3, full coverage of the CAEX vocabulary according to R2 can be achieved. Since the new concepts cover the changes introduced by the new AutomationML version, the resulting ontology also covers Requirement R4.
AttributeTypeLibrary and AttributeTypes
The recent AutomationML-Standard IEC 62714-2[14] introduced both AttributeTypes and AttributeTypeLibraries. These can be used to give Attributes of e.g. InternalElements a consistent and reusable semantic description. They are structured in the same way as e.g. the InterfaceClasses and InterfaceClassLibraries. Therefore, in order to allow the modeling of AttributeTypes in the ontology, two classes were added to the AutomationML-Ontology: AttributeType, which allows the modeling of a template that can be reused for Attributes and AttributeTypeLib, which contains and defines AttributeTypes. These classes are linked to the remaining concepts of the ontology through newly defined object properties: The object property hasRefAttributeType is a direct translation of the CAEX-Tag RefAttributeType. In AutomationML it serves two purposes: First, it links an Attribute to the AttributeType it is based on. Second, it can also link an AttributeType to the AttributeType it is derived from.
Constraints
In AutomationML, the values of attributes can be restricted using three different types of constraints:
A NominalScaledType-constraint restricts the values of an attribute to a set of allowed values, while an OrdinalScaledType-constraint restricts the values to a range specified by a minimal and a maximal attribute type. Additionally, an UnknownType-constraint exists that can be used for any other type of constraint.
In order to allow the use of these constraints in the AutomationML-Ontology, the classes NominalScaledConstraint, OrdinalScaledConstraint and UnknownTypeConstraint were added to the ontology, alongside data properties to define their constraints.
MappingObjects
In AutomationML, MappingObjects map the attributes or interfaces of an InternalElement to the corresponding attributes or interfaces defined in RoleClasses or SystemUnitClasses. If the names of the Attributes or Interfaces match, this is done implicitly, otherwise an explicit MappingObject is created. This ensures that the data and functionalities defined in the InternalElement are correctly associated with the corresponding elements in the RoleClass/SystemUnitClass. This allows for clear and consistent interpretation of the roles and properties within the AutomationML model.
4.3 Context dependent Vocabulary
On top of the CAEX vocabulary, AutomationML uses regular CAEX concepts which can have different meanings depending of their context. To end users, these concepts are only ever shown in their contextual meaning (e.g. as a Facet instead of an InternalElement). In order to increase usability for the end user, these relationships should be made explicit in the ontology in accordance with Requirement R3.
MirrorObjects
In AutomationML, a MirrorObject is used to reference and replicate elements from one part of the model into another without duplicating the actual data. It creates a ”mirror” of an existing object, such as an InternalElement, SystemUnitClass, or other elements, allowing the same object to be used in multiple locations within the model.
The MirrorObject references the original object by its ID, ensuring that any changes to the original are automatically reflected in all mirrored instances. This promotes consistency and reduces redundancy within the AutomationML model by reusing existing data structures instead of copying them. MirrorObjects are not part of the CAEX vocabulary, but can instead be interpreted based on their context. In order to be able to explicitly model Mirror Object relationships, the object properties hasMirrorObject and hasMasterObject have been added to the ontology. These properties connect InternalElements or Interfaces with their respective Mirror/Master-Objects.
Facet
A Facet is a mechanism used to group related attributes, interfaces, and sub-elements within a single object. It allows for the modular and reusable definition of properties and behaviors, which can then be applied to various elements in the model, enhancing consistency and simplifying the management of complex systems. Facets help in organizing and structuring the data more efficiently by encapsulating related aspects of a model element. While Facets are not part of the CAEX vocabulary, they are commonly used in AutomationML to create domain specific views of a component and are therefore explicitly included as a class in the ontology.
Group
In AutomationML, a Group is a mechanism used to logically organize and manage multiple elements, such as InternalElements or SystemUnitClasses, into a collective structure. Groups allow for the aggregation of related elements, facilitating easier handling, reuse, and understanding of complex models within a project. Like Facets, they are not part of the CAEX-vocabulary, but are explicitly included in the ontology as a class.
4.4 Design Practices and Reuse
The AutomationML ontology is published under W3ID, which is a permanent identifier, that is administered by the W3C Permanent Identifier Community Group. It is permanently available using the ontology base IRI w3id.org/hsu-aut/AutomationML. Administration and maintenance of the ontology is supported by the Institute of Automation Technology at Helmut-Schmidt-University. Both the ontology itself as well as its classes and relationships are annotated with metadata documenting their creation process and correct usage.
4.5 Ontology Summary
The AutomationML-Ontology contains a total of 22 classes, 8 ObjectProperties and 5 Data Properties. Their terminology remains very close to the terms used in AutomationML (R1). Together, these classes and properties allow the full coverage of both the CAEX-Vocabulary according to Requirement R2 as well as the AutomationML-specific vocabulary according to Requirement R3. By updating the terms to cover all newly added terms from the most recent specification, it allows full coverage of the current version of AutomationML (R4).
5 Mapping AutomationML to OWL
While Section 4 assured that the AutomationML-Ontology provides a comprehensive vocabulary to cover the entirety of the AutomationML-content, the following Section provides:
1.
An alignment of AutomationML vocabulary with OWL-specific vocabulary (e.g. rdf:type etc.),
2.
A formal description of the relationships between CAEX, AutomationML, and OWL vocabulary,
3.
A formal description of context dependent AutomationML concepts,
The mapping rules in this section - as well as all other rules needed to automatically populate an AutomationML ontology with the data from an AutomationML file - are available in the RDF-Mapping-Language RML.
5.1 Creation of IRIs
Every resource inside an RDF graph has to be identified by an IRI, that distinguishes it from other resources. For AutomationML files, two main ways exist to create these identifiers: The first option relies on the IDs that are provided in the AutomationML file. While this approach is straightforward, IDs are only mandatory for InternalElements and ExternalInterfaces, where AutomationML also uses these IDs internally for references. For all other elements in AutomationML files, the use of IDs is optional and can therefore not reliably be used for referencing. Instead, elements without an ID are referenced through a reference path, which is the second option to create unique IRIs. In AutomationML, these elements must therefore have a name that is unique to their level in the element hierarchy.
In the RML mapping, this reference approach was mimicked: IDs are used as IRIs where they are mandatory (i.e. for InternalElements and ExternalInterfaces), while IRI-safe ReferencePaths are used for all remaining elements. References to AttributeTypes, Role-, SystemUnit- and InterfaceClasses take the form of reference paths as well, which makes the creation of ObjectProperties between instances and their respective classes straightforward. All instances are further given non-unique human-readable labels (via rdfs:label) based on their names in AutomationML.
5.2 Creation of Classes and Instances
For the mapping process, a decision had to be made which AutomationML concepts would be mapped as classes and as instances. For InternalElements and ExternalInterfaces, this decision is trivial: They are unique individuals, as evidenced by the mandatory use of an ID.
Conceptually, classes in AutomationML are defined slightly different from classes in OWL. On one hand, they function like OWL classes - i.e. as a conceptual definition of a group of objects. On the other hand, classes in AutomationML can also function as a template for instances - i.e. they can posses attributes or be composed of multiple subcomponents. In this sense they show characteristics of both OWL classes as well as OWL instances.
OWL 2 DL introduced the concept of punning[24]. It allows repurposing the same identifier to represent both a class and an instance. Using this mechanism, AutomationML classes can be represented as OWL classes as well as OWL instances simultaneously. Punning is therefore used extensively in the mapping of RoleClasses, SystemUnitClasses, InterfaceClasses and AttributeTypes.
In the class view, any RoleClass, SystemUnitClass, InterfaceClass, or AttributeType is converted to an OWL class. Class membership is implied, if instances reference the class using hasRefBaseSystemUnitClass, hasRefBaseClass, or hasRefAttributeType respectively (s. Equation 4).
In the instance view, any RoleClass, SystemUnitClass, InterfaceClass or AttributeType can be further described with any additional information from the AutomationML file. In this context, the OWL classes SystemUnitClass, RoleClass, InterfaceClass, and AttributeType are be treated as metaclasses, that define the structure and behavior of other classes. Classes defined in AutomationML are linked to their respective metaclasses using rdf:type.
5.3 Mapping of Roles
As described in Section 4, RoleClasses define the abstract functions of either InternalElements or SystemUnitClasses. AutomationML allows two types of RoleClass assignments:
RoleRequirement can be used to describe roles that an InternalElement or a SystemUnitClass is required to fill in a plant.
SupportedRoleClass can be used if an InternalElement or a SystemUnitClass is capable of filling a role.
These relationships are modeled in two complementary mechanisms: If an InternalElement or SystemUnitClass references a RoleClass via SupportedRoleClass, it is considered to be an instance of that class (s. Equation 5). Additionally, the object properties hasRoleRequirement and hasSupportedRoleClass are assigned to explicitly distinguish between required and supported RoleClasses.
5.4 Mapping of Class Hierarchies
AutomationML allows modeling class hierarchies, in which subclasses can be derived from base classes. Inheritance between classes has to be explicitly modeled using RefAttributeType for AttributeTypes or RefBaseClassPath for SystemUnit-, Interface-, and RoleClasses.
In the SystemUnitClass-, RoleClass-, InterfaceClass-, and AttributeTypeLibraries, this inheritance mechanism results in a class hierarchy using rdfs:subClassOf according to Equations 6 and 7.
5.5 Mapping of Internal Structures
Some elements in AutomationML can have an internal structure, i.e. they consist of multiple individual components. For example, both InternalElements and SystemUnitClasses can contain other InternalElements, Interfaces, and Attributes, that can, in turn, have an internal structure as well. In the same way, ExternalInterfaces can contain other ExternalInterfaces (e.g. a voltage interface on USB) and Attributes can contain other Attributes (e.g. x and y values on coordinates). In AutomationML, this relationship can be inferred from the XML-hierarchy in the source document. In the ontology, this implies a PartOf-Relationship. It is therefore modeled using the object property hasPart.
5.6 InternalLinks
In AutomationML, an InternalLink describes a directional connection between two interfaces. InternalLinks are crucial because they establish relationships between different components within a system model. The mapping uses this information in two ways: On one hand it follows the AutomationML terminology, in which an InternalLink instance has two RefPartnerSides. However, in comparison to AutomationML, OWL offers more suitable ways of representing links between elements. Additionally, an InternalLink is therefore represented using the :isLinked-Property that links both RefPartnerSides of an InternalLink while maintaining directionality.
5.7 Mirror-Objects
Depending on the context, the RefBaseSystemUnitPath is used in AutomationML to describe two different relationships. If it connects an InternalElement with a SystemUnitClass, the InternalElement is based on the SystemUnitClass (s. Section 5.2). However, if it connects an InternalElement with another InternalElement, it implies a Mirror-Relationship between the two elements. In this case, the second element follows all changes made to the first one. By introducing the :hasMirrorObject and :hasMasterObject, this connection is made explicit in the ontology.
5.8 Mapping-Objects
In AutomationML, MappingObjects are used to link and correlate elements from different models, ensuring interoperability and data consistency across various engineering domains and tools. If e.g. an InternalElement and its RoleClass have corresponding Attributes or Interfaces with matching names, a MappingObject is implicitly assumed. If the names do not match, the relationship is explicitly stated in the AutomationML file. In the latter case, the mapping is straightforward. In the former case, the mapping information has to be extracted from the data according to Equations 10 / 11.
5.9 Facets and Groups
Both Facets and Groups are not explicitly part of the original CAEX-Vocabulary. Instead, they are modeled as InternalElements that have a RoleRequirement of Group or Facet from the AutomationMLBaseRoleClassLibrary. Their membership in the class Group/Facet is therefore inferred according to Equations 12 and 13.
5.10 Inheritance of Subcomponents
In AutomationML, the internal structure of a class is inherited from a parent class to the child class. In the AutomationML model, this relationship is not explicitly modeled, but can instead be inferred through the inheritance relationship. It is additionally possible to override inherited objects.
For the mapping to RDF, this means that an internal class structure is inherited from the parent class if the inherited structure was not overridden (s. Equation 14).
5.11 Mapping Process and Usage
Fully mapping an AutomationML file to RDF triples is a two-step process. In the first step, the provided RML mapping is executed. This populates the knowledge graph with all RDF triples corresponding to the contents of the file. In a second step, a set of SPARQL-Queries is executed, which adds additional triples, that cannot otherwise be created by means of pure RML mapping. This information includes the inheritance of Attributes and Interfaces from a parent element, as well as creating explicit MappingObjects of Attributes and Interfaces.
A mapping application is available***https://github.com/hsu-aut/aml2owl, that allows automatic mapping of a file from the command line. It includes both the RML mapping as well as the SPARQL queries.
Alternatively, the mapping and the SPARQL queries are individually available on GitHub†††https://github.com/tomwestermann/AutomationML2Ontology. The mapping can be executed using any RML-compliant mapping tool, while the SPARQL queries can be executed using any compliant SPARQL processing tool.
6 Mapping Example
To illustrate the mapping rules from the previous section, a Four-Tank Mixing System will be used as a running example. Further models are available on Github.
6.1 System Description: Four-Tank Mixing System
In the main use case, a laboratory scale process plant in the Institute of Automation Technology at Helmut-Schmidt-University is considered. It is used for fluid mixing and consists of three dosing tanks, a mixing tank, eleven valves and two pumps. Pipes connect the various components for fluid transport. A Piping and Instrumentation Diagram (P&ID) of the system can be seen in Figure 3. The system’s topology was described in AutomationML according to the best practice recommendations by the AutomationML society.
The remainder of this section shows how the main concepts of AutomationML are translated to RDF-Triples using excerpts from the AutomationML file of this system.
6.2 Creation of IRIs
As described in Section 5.1, the creation of IRIs for the ontology follows the reference mechanisms that AutomationML is using internally. Therefore, IDs are only used for the IRI creation of InternalElements and ExternalInterface. For all other elements, unique paths are created using the ”Name”-Attribute. This mimics AutomationMLs approach to references.
Listing 6.2 and Figure 4 show an example of this logic.
While the InternalElement B201 is assigned an IRI from its ID, both the Attribute Length as well as the AttributeType Length use their paths as an IRI.
{listing}[ht]
Example of an InternalElement with an Attribute ”Length”.
{Verbatim}[commandchars=
¡InstanceHierarchy Name=”MyIH”¿
¡InternalElement Name=”B201”
ID=”44806a23-d2bd-45d2-8344”¿
¡Attribute Name=”Length”
RefAttributeType=”MyAtLib/Length”¿
¡/Attribute¿
¡/InternalElement¿
¡/InstanceHierarchy¿
6.3 Creation of Classes and Instances
The following section provides an example of the creation of classes and instances. Listing 6.3 shows an example of an InternalElement, that has all major elements that AutomationML uses: A SystemUnitClass, an Attribute, an ExternalInterface as well as two RoleClasses - one as a SupportedRoleClass and one as a RoleRequirement.
[ht]
InternalElement with all basic AutomationML elements.
{Verbatim}[commandchars=
¡InternalElement Name=”B201”
ID=”44806a23-d2bd-45d2-8344”
RefBaseSystemUnitPath=”MySucLib/Vessel”¿
¡Attribute Name=”Length”
RefAttributeType=”MyAtLib/Length”¿
¡/Attribute¿
¡ExternalInterface Name=”Input”
ID=”ce45d7b3-169d-4be8-9746”
RefBaseClassPath=”MyIcLib/Port”¿
¡/ExternalInterface¿
¡SupportedRoleClass
RefRoleClassPath=”MyRcLib/Vessel” /¿
¡RoleRequirements
RefBaseRoleClassPath=”MyRcLib/DosingTank” /¿
¡/RoleRequirements¿
¡/InternalElement¿
Figure 5 shows the corresponding relationship between OWL instances and classes from Listing 6.3: For example, the InternalElement B201 is:
an instance of the SystemUnitClass Vessel, since it is its RefBaseSystemUnitClass,
an instance of the RoleClass Vessels, since it supports that RoleClass,
not an instance of the RoleClass DosingTank, since having a RoleRequirement does not imply class membership.
The Attribute Length and the Interface Input are also instances of their respective AttributeTypes and BaseClasses. The RDF representation of the AutomationML data additionally connects the instances to their classes using the AutomationML terminology.
6.4 Mapping of Class Hierarchies
The previous section showed how an InternalElement (or its Attributes and Interfaces) can relate to OWL classes that are derived from AutomationML libraries.
Since AutomationML allows inheritance, these OWL classes become part of a class hierarchy. AutomationML additionally provides further meta information about these classes (e.g. default values, semantic references, provenance, etc.). To adequately capture this information, punning, i.e. the simultaneous use of the same IRI as a class and an instance, is used.
The example in Listing 6.4 and Figure 6 shows how the AttributeType Length from the previous example is mapped and embedded into an OWL class hierarchy. For AttributeTypes, the rdfs:subClassOf-relationship can be inferred from the RefAttributeType in AutomationML.
[ht]
Excerpt from the AttributeLibrary.
{Verbatim}[commandchars=
¡AttributeTypeLib Name=”MyAtLib”¿
¡AttributeType
Name=”Dimensions”¿
¡AttributeType Name=”Length”
RefAttributeType=”MyAtLib/Dimensions” /¿
¡/AttributeType¿
¡/AttributeTypeLib¿
OWL Class hierarchies are created from all AutomationML libraries, i.e. also from the RoleClass-, SystemUnitClass-, and InterfaceClassLibraries. The mapping approach is similar for all other libraries.
6.5 Mapping of InternalLinks
Listing 6.5 shows an example of two InternalElements, whose interfaces are connected by an InternalLink. Figure 7 shows both mechanisms of linking their Interfaces. On one hand, they are connected in a way that very closely follows the AutomationML-vocabulary with an instance of the class InternalLink that connects the Interfaces through the Object Property :hasRefPartnerSideA/B. On the other hand, they are connected directly through the object property :isLinked. This makes traversing the connections between InternalElements (e.g. for queries) more intuitive, while preserving the terminology most familiar to AutomationML practitioners.
[ht]
Example of an InternalLink that connects a Pipe with a Tank.
{Verbatim}[commandchars=
¡InstanceHierarchy Name=”MyIH”¿
¡InternalElement Name=”Pipe”
ID=”a20e3eac-9ec0-41f1-852a”¿
¡ExternalInterface Name=”Output”
ID=”6eea7a40-43fd-4aee-a1d3-”¿
¡/ExternalInterface¿
¡InternalLink Name=”Pipe˙B201”
RefPartnerSideA=”6eea7a40-43fd-4aee-a1d3”
RefPartnerSideB=”ce45d7b3-169d-4be8-9746”/¿
¡/InternalElement¿
¡InternalElement Name=”B201”
ID=”44806a23-d2bd-45d2-8344”¿
¡ExternalInterface Name=”Input”
ID=”ce45d7b3-169d-4be8-9746”¿
¡/ExternalInterface¿
¡/InstanceHierarchy¿
7 Usage
A reliable and reproducible way of translating AutomationML-Files to RDF-Triples opens up a plethora of new applications using the different technologies from the semantic web stack (s. Section 2.3). In this section, two such applications, namely querying and validation, and their advantages will be described. The first application describes the use of the query language SPARQL to construct complex graph queries. The capabilities of SPARQL are compared to xPath and xQuery, which are considered the most common query languages that operate directly on the XML representation of AutomationML. The second application describes an automatic validation of an AutomationML file by using SHACL to formulate validation patterns. SHACL’s capabilities are then compared to XSD and Schematron, two common approaches to validate the content of XML files.
7.1 Complex Graph Queries
As stated in Section 2.2, the AutomationML model describes a labeled and directed multigraph, that is serialized as an XML file in the shape of a Tree.
Common tools to perform queries on XML data (like xPath or xQuery) were designed to navigate tree structures, and are therefore lacking mechanisms to handle the additional edges in the graph created by ReferencePaths and InternalLinks.
However, many industrial applications could significantly benefit from the possibility to query the full spectrum of information provided by the AutomationML file.
To illustrate the advantages of such graph queries, we revisit the Four-Tank-Mixing System described in Section 6.1. We will compare the extraction of information using different approaches: xPath and xQuery, which operate directly on the XML logic underlying the AutomationML file, and SPARQL, which queries an RDF graph generated from the AutomationML file.
Sets of search queries of varying complexity will be considered to highlight the strengths and specific limitations of each approach, demonstrating the potential benefits of employing graph queries in relevant scenarios.
Use Case 1: Conditional Element-Selection
In the first Use Case, the simple selection of elements based on some selection criteria was considered. Three different scenarios were designed. Each scenario differs in the degree of separation in the transformed graph between the element that should be selected and the element at which the selection criteria is applied.
1.
Scenario 1 - First Degree Selection:
Select all InternalElements A that have a certain RoleRequirement B:
2.
Scenario 2 - Second Degree Selection:
Select all InternalElements A that have a RoleRequirement B, where B references a RoleClass C as a RefBaseClass:
3.
Scenario 3 - nth Degree Selection:
Select all InternalElements A that have a RoleRequirement B, where B references a RoleClass C as a RefBaseClass through transitive closure.
Example queries of these three scenarios were applied to the the Mixing System: All InternalElements should be selected that have the RoleRequirement of being 1) a Vessel, 2) a Plant Equipment (or any direct descendant) or 3) a Resource (or any transitive descendant). For the example in Figure 8, the first scenario should return just B201, while the queries from the other scenarios should return both InternalElements.
For all three of these Scenarios, queries were written in each of the three query languages.
[ht]
xPath, xQuery and SPARQL Query for Scenario 1.
xPath:
{Verbatim}[commandchars=
//InternalElement[RoleRequirements[
@RefBaseRoleClassPath=’MyRcLib/Vessel’]]
xQuery:
{Verbatim}[commandchars=
for $ie in //InternalElement
where $ie/RoleRequirements/
@RefBaseRoleClassPath = ”MyRcLib/Vessel”
return $ie
SPARQL:
{Verbatim}[commandchars=
SELECT ?ie
WHERE –
?ie rdf:type aml:InternalElement .
?ie aml:hasRoleRequirement aml:MyRcLib/Vessel.˝
The queries regarding the first scenario were easily achievable in all three languages (s. Listing 8).
However, both the second and third scenario make use of the AutomationML reference paths. The query in Scenario 2 not only references the RoleRequirement of InternalElement A, but also the RoleRequirement’s RefBaseClass. In the XML tree, this information is located at a different branch of the tree at the RoleRequirement node and thus necessitates a traversal of the graph that does not follow the hierarchical structure of the XML file. Scenario 3 further extends this by combining the filter condition with an open-ended transitive closure statement, i.e. a sequence of RefBaseClassPaths of indeterminate length.
[ht]
xQuery for Scenario 2 and SPARQL Query for Scenario 3.
xQuery:
{Verbatim}[commandchars=
let $roleClasses :=
//RoleClass[@RefBaseClassPath=’PlantEquipment’]
for $ie in //InternalElement
where some $rc in $roleClasses satisfies
$ie/RoleRequirements/@RefBaseRoleClassPath =
string-join($rc/ancestor-or-self::*/@Name,’/’)
return $ie
SPARQL:
{Verbatim}[commandchars=
SELECT ?ie
WHERE –
?ie rdf:type aml:InternalElement .
?ie aml:hasRoleRequirement/(rdfs:subClassOf)*
aml:PlantEquipment .˝
Both xPath and xQuery operate directly on the XML-tree, and have therefore no mechanisms to handle ReferencePaths. xQuery offers some mechanisms to circumvent this limitation: Technically, the second scenario could still be solved using xQuery. The approach includes reconstructing the reference path of all RoleClasses, iterating over them, and individually comparing them to the desired value (s. Listing 8). However, this approach is very limited in scope, e.g. capturing the open-ended transitive closure statement from Query 3 in this way is not possible. Using SPARQL, due to the support of Property Paths, all three queries can be answered in a much more compact and arguably more intuitive form (s. Listing 8).
Use Case 2: Automatic Model Creation
In the second use case, a more complex query will be formulated, with the aim of automatically producing a Material Flow Graph of the Four-Tank-Mixing System. The automatic generation of material flow graphs or other causal graphs has many applications in areas like the automatic generation of simulation models[25][26], or the contextualization and diagnosis of data anomalies[27] and alarm floods[28][29][30].
As in the previous Use Case, two scenarios with different levels of complexity will be covered:
1.
Scenario 1 - Bidirectional FlowGraph:
Create a flow graph assuming bidirectional flow.
2.
Scenario 2 - Valve-state dependent FlowGraph:
Create a valve-state dependent directional flow graph.
In the first scenario, the content of the AutomationML-file should be used to create a material flow graph of all possible paths through the system. In the second scenario, the query was extended to include information about the state of valves, as they enable or disable product flow in parts of the system.
Therefore, both queries build upon the same basic graph pattern, but extend it with additional conditions that have to be met.
In this use case, it is assumed that the material flow interface in AutomationML was modeled using the InterfaceClass Port or interfaces that inherit from this class. The Port-InferfaceClass is defined in the AutomationMLBaseInterfaceLibrary. In addition, any valves in the model can have a boolean ValveState-Attribute that indicates, whether the valve is open or closed.
Since it does not support multi-step-patterns, xPath was unable to create a material flow graph for either of the three scenarios. Using xQuery, a Flow Graph for the first Scenario could be created. However, adding additional checks and constraints quickly results in very deeply nested loops over all possible connections. While technically possible, the queries this approach produced were opaque and error-prone.
SPARQL, on the other hand, offers the possibility to define repeating patterns inside the graph that should be traversed. Due to these features, all queries can be answered using a more straightforward and compact SPARQL-Query.
A SPARQL query that solves the second scenario can be found in Listing 9.
[ht]
SPARQL query, that creates a material flow graph while considering valve states.
{Verbatim}[commandchars=
construct–?s aml:flows ?o.˝ where –
?s a aml:InternalElement.
?o a aml:InternalElement.
?s aml:hasInterface ?i1.
?o aml:hasInterface ?i2.
?i1 aml:isLinkedTo ?i2.
OPTIONAL –
?s aml:hasAttribute ?vs.
?vs aml:hasRefAttributeType aml:ValveState.
?vs aml:hasAttributeValue ?state.
Filter(!bound(?state) —— ?state = ’true’).
Comparison of xPath, xQuery, SPARQL
The three query languages show differences in their ability to extract information from the AutomationML model. All three query languages are able to perform basic queries that encompass selecting elements based on filtering conditions. However, once the AutomationML-ReferencePaths become involved, difficulties emerge in xPath and xQuery.
For this comparison, we refer to the different categories of filtering conditions from Section 7.1: In First Degree Filtering and Comparisons, the filter condition applies to the direct attributes that an element has. These tasks also encompass all queries in which only the ReferencePath is of interest, and the information at the destination of the path is irrelevant. All query languages were able to fulfill these tasks. In Second-Order Filtering, the filter condition applies to an element that the first element references via a reference path. For these queries, xPath is unable to retrieve the targeted elements. xQuery, on the other hand, is technically able to answer those queries through a combination of loops and string reconstructions and comparisons. However, the resulting queries are complicated and not robust.
In n-th order filtering, the filter condition applies to an element that the first element references through a chain of two or more subsequent reference paths. This case also includes use cases like multilevel inheritance or complex path patterns. In these cases, xQuery is not suitable to create functional queries. All of these shortcomings of xPath and xQuery to query AutomationML files stem from the mismatch between the XML tree structure that both languages were designed for, and the labeled directed multigraph that the contents of an AutomationML file actually represent. Since it is a dedicated graph query language, SPARQL is able to handle such use cases, while also offering explicit support for more advanced and graph specific features like property paths and complex graph patterns.
In summary, the comparison between the query languages shows that xPath and xQuery are at best sufficient for simple queries and information retrieval. If the scenarios become more complex, SPARQL offers greater functionalities and ease of use.
7.2 Data Validation
Besides the potential for querying, representing an AutomationML file in the form of an ontology also offers possibilities for validation of the file’s content. A case study was designed to compare three common validation languages - XSD, Schematron and SHACL - for their suitability for AutomationML data.
XML Schema, usually abbreviated XSD, operates directly on the XML files. It is commonly used to validate the syntactic correctness of an XML file. Schematron is an XML schema language used for rule-based validation of XML documents. In comparison to XSD, Schematron offers a more flexible, rule-based approach. SHACL on the other hand is a language for validating RDF data. It provides mechanisms to define constraints on RDF graphs, ensuring that data conforms to specific shapes and patterns.
Use Case: Library-Level Rules
In this use case, the validation concerns the correct usage of elements that were imported from user defined libraries. These libraries are commonly accompanied by textual specifications that define the proper usage of the imported elements. As those specifications lack a formal or machine interpretable definition, their correct usage needs to be manually checked by domain experts. For this use case, the Application Recommendation Automation Project Configuration (AR APC) was considered. It is a framework that describes guidelines for the configuration of automation projects.[31]
For this example, the SensorPort-RoleClass was considered (s. Figure 11). It is a RoleClass used to model the connection between devices for the transfer of sensor signals. As such, it can contain an Interface that connects to SensorPortsInterfaces. The AR APC defines two constraints on this class:
Scenario 1 - Cardinality Constraint:
All Instances of the class can have at most one SensorPortInterface.
Scenario 2 - Connection Constraint:
SensorPort-Interfaces can only be connected to other SensorPort-Interfaces.
[ht]
Schematron and SHACL constraints for Scenario 2.
Schematron:
{Verbatim}[commandchars=
¡rule context=”caex:InternalElement[
caex:RoleRequirements[@RefBaseRoleClassPath=
’ArApcRcLib/SensorPort’]]”¿
¡assert test=”
every $link in caex:InternalLink satisfies
(//caex:ExternalInterface[
@ID=$link/@RefPartnerSideA]/@RefBaseClassPath=
’ArApcIcLib/SensorPortInterface’ and
//caex:ExternalInterface[
@ID=$link/@RefPartnerSideB]/@RefBaseClassPath=
’ArApcIcLib/SensorPortInterface’)”¿
¡/assert¿
¡/rule¿
SHACL:
{Verbatim}[commandchars=
:SPIConnectionShape
a sh:NodeShape ;
sh:targetClass aml:SensorPortInterface;
sh:property [
sh:path aml:isLinkedTo;
sh:class aml:SensorPortInterface;] .
Comparison of XSD, Schematron and SHACL
Both scenarios constitute conditional constraints, i.e. they check the value of an XML-Attribute (RefBaseClassPath) and only apply the validation if it matches a certain value (CommunicationRoleClassLib/PhysicalDevice/SensorPort).
While XSD can validate basic data types and structure, it doesn’t support conditional validation and thus lacks the ability to handle complex relationships between elements or logic. Compared to XSD, Schematron and SHACL are far more expressive and flexible. However, Schematron is tied to XML’s tree structure and can’t natively handle ReferencePaths or InternalLinks. Solving validation scenarios thus involves complex iterations over the data, just as it was the case with xQuery.
Between the three approaches, XSD is most suitable for basic structural validation (e.g., ensuring elements appear in the correct order or enforcing data types), while both Schematron and SHACL can be used for cross-element validation with a lower level of complexity (s. Table 4). Compared to the other validation approaches, SHACL has access to a semantically described relationship model and can therefore define complex graph shapes that the data must conform to. As the complexity of the validation pattern increases, SHACL’s advantages are expected to grow more pronounced.
Validation Property | XSD | Schematron | SHACL |
Basic Structural Validation | + | + | + |
Library-Level Cardinality Validation | - | + | + |
Library-Level Connection Constraint | - | o | + |
8 Conclusion
AutomationML is a widespread and expressive exchange format in the automation domain. In the past, multiple attempts have been made to incorporate the knowledge contained in AutomationML files into knowledge graphs. While these approaches have been partially successful, no common method of mapping the information contained in AutomationML-files to knowledge graphs exists. To mitigate this issue, this article provides:
1.
An AutomationML ontology that is compatible with version 3 of the CAEX specification.
2.
A declarative mapping that offers practitioners a straightforward transformation of AutomationML files to RDF triples.
The ontology‡‡‡https://github.com/hsu-aut/IndustrialStandard-ODP-AutomationML as well as the mapping files and applications§§§https://github.com/hsu-aut/aml2owl are freely available to practitioners and can be found on Github. They were validated on multiple exemplary AutomationML files and were able to capture and adequately replicate the information contained in the file.
Two studies were conducted to evaluate the mappings’ potential for graph queries and validation. They showed that the mapping to RDF offers expanded possibilities for querying as well as data validation.
8.1 Discussion
In comparison to previous approaches to mapping AutomationML files to RDF graphs, the use of an RML mapping offers multiple advantages. For one, all used artifacts (RML mapping and SPARQL queries) are independent of any implementation language. In case of updates to AutomationML specification, only the mapping needs to be updated, while the tools that use them can remain the same. This should result in lower total maintenance. The evaluation showed that mapping AutomationML files to RDF enables more complex use cases for querying and validation. Since common XML-based solutions were designed to handle a different type of graph structure, they can only operate on a subset of the information that AutomationML actually provides.
The experiments regarding automatic validation show that mapping the AutomationML file to RDF and subsequently applying SHACL validation rules improve upon the validation capabilities of the current approach using XSD. The same conclusion was reached regarding the querying capabilities, in which SPARQL proved to be more capable and user friendly than xPath and xQuery.
8.2 Outlook and Future Work
This work is part of an ongoing effort to create a formal description of AutomationML. A plug-in for the AutomationML Editor is currently in development. It can be used to translate an AutomationML file to RDF and then query it using SPARQL or evaluate it against SHACL validation rules.
The artifacts provided by this work enable new uses of the information in an AutomationML file. The capability to formulate complex queries and create models based on the file contents allow new applications in the field of informed machine learning, in which prior knowledge of a system is used to improve various common AI tasks such as anomaly detection or diagnosis. These approaches would especially benefit from a connection between the static information from the AutomationML file with run time data from the system. Additionally, having a simple way to represent the files contents in an RDF graph allows the integration of the knowledge from the AutomationML file with other information artifacts about a system.
Acknowledgment
We would like to express our gratitude to the members of the FD4AML working group from the AutomationML association for their valuable feedback and constructive criticism, which significantly contributed to the improvement of this work. In particular, we thank Miriam Schleipen, Tina Mersch, Sandra Zimmermann and Aljosha Köcher for their insightful comments and suggestions that helped refine the analysis and presentation of our findings.